OpenStudy (anonymous):

What is the solution to the equation 9(w – 4) – 7w = 5(3w – 2)?

OpenStudy (helder_edwin):

first apply the distributive law and then cancel out the corresponding terms

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

\[ 9(w-4)-7w=5(3w-2) \] \[ 9w-36-7w=15w-10 \] u get this, right?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

and yeah